# FlowAlign: Trajectory-Regularized, Inversion-Free Flow-based Image Editing

### v2

@@ Abstract
Recent inversion-free, flow-based image editing methods such as FlowEdit leverages a pre-trained noise-to-image flow model such as Stable Diffusion 3, enabling text-driven manipulation by solving an ordinary differential equation (ODE). While the lack of exact latent inversion is a core advantage of these methods, it often results in unstable editing trajectories and poor source consistency. To address this limitation, we propose FlowAlign, a novel inversion-free flow-based framework for consistent image editing with principled trajectory control. FlowAlign introduces a flow-matching loss as a regularization mechanism to promote smoother and more stable trajectories during the editing process. Notably, the flow-matching loss is shown to explicitly balance semantic alignment with the edit prompt and structural consistency with the source image along the trajectory. Furthermore, FlowAlign naturally supports reverse editing by simply reversing the ODE trajectory, highlighting the reversible and consistent nature of the transformation. Extensive experiments demonstrate that FlowAlign outperforms existing methods in both source preservation and editing controllability.
@@ 1 Introduction
In text-based image editing Meng et al. (2021); Mokady et al. (2023); Tumanyan et al. (2023); Hertz et al. (2023); Kim et al. (2024); Kulikov et al. (2024); Brooks et al. (2023), the goal is to transform a source image into a target based on either textual descriptions of the images or specific editing instructions. From a distributional perspective, the image editing task can be interpreted as a continuous normalizing flow (CNF) Papamakarios et al. (2021) that pushes forward a source distribution to a target distribution. Specifically, we parameterize a velocity field, that uniquely determines the flow, using a neural network. Then, the generative process corresponds to solving an ordinary differential equation (ODE) governed by the trained velocity field. To reduce the computational cost of simulating ODEs during training for likelihood evaluation, flow matching has been proposed Lipman et al. (2023). Its conditional variant enables direct supervision by computing the target velocity, allowing efficient training of flow models between arbitrary distributions. These flow models includes score-based diffusion models Song et al. (2021b); Ho et al. (2020); Song et al. (2021a) as well as rectified flow models Liu et al. (2023b); Esser et al. (2024).
@@
Recently, within the framework of flow-based models, text-to-image generation has achieved significant advancements through improved time discretization, loss weighting, and model architecture, notably based on DiT Peebles and Xie (2023). While these foundational models are trained to map samples from a normal distribution to a clean data distribution, additional training—either from scratch or via fine-tuning—is necessary to establish a flow between two arbitrary distributions.
@@
To mitigate this additional cost for constructing a new flow between image distributions, several approaches such as SDEdit Meng et al. (2021) and Dual Diffusion Implicit Bridge (DDIB) Su et al. (2023) have leveraged pre-trained noise-to-image diffusion models. However, SDEdit requires careful selection of an appropriate initial noise level for editing, while DDIB relies on an inversion process that is prone to errors arising from discretization and an approximated velocity field for subsequent timesteps. Recently, RF-inversion Rout et al. (2025a) proposes an optimal-control based inversion for flow models, in which a guiding vector field steers the dynamics toward high-likelihood samples via velocity interpolation. Unfortunately, it still involves computational overhead due to the ODE inversion. To address this, FlowEdit Kulikov et al. (2024) proposed simulating an ODE between two image samples without inversion. However, empirical results showed that the method is quite sensitive to the hyperparameters and often fails to retain the source consistency. Moreover, the method heuristically applies classifier-free-guidance (CFG) to the both source and target velocity fields with different scaling factor and relies on skipping early timesteps of the ODE to enhance editing quality. These design choices introduce multiple hyperparameters to be searched and compromise the deterministic nature of ODE (see Section 6).
@@
Motivated by the observation that these limitations of FlowEdit arise from nonsmooth and unstable editing trajectories—partly due to the lack of explicit latent inversion—we introduce FlowAlign, a flow-based inversion-free approach for consistent and controllable text-driven image editing through trajectory regularization. Specifically, to overcome the instability caused by the absence of the inverted latent, we cast the editing process as an optimal control problem, where the deviation from the smooth trajectory is incorporated as a flow-matching penalty term. This formulation also allows for direct simulation between the source and target images through an ODE. Although the regularization is primarily designed to encourage geometrically meaningful trajectories, we further observe that it also naturally enforces source consistency by penalizing unnecessary deviations from the original image. This dual effect enables the editing trajectory to maintain a robust balance between semantic alignment and structural fidelity.
@@
Our method is also computationally efficient, requiring only one additional function evaluation (NFE) per ODE step for the regularization term. In contrast, methods like FlowEdit Kulikov et al. (2024) incur twice the NFE due to the reliance on classifier-free guidance. Despite this efficiency, FlowAlign achieves superior source preservation and competitive or improved editing quality compared to existing approaches in image, video, and 3D editing. Finally, FlowAlign supports editing through backward ODE with the learned flow field, enabling accurate reconstruction of the original image from the edited output. This highlights the reversible and deterministic nature of the learned transformation, made possible by our explicit trajectory regularization.
@@ 2 Backgrounds
Suppose that we have access to samples from two distributions $X_{1}\sim p$ and $X_{0}\sim q$ that forms independent coupling, $\pi_{0,1}(X_{0},X_{1})=p(X_{1})q(X_{0})$. We can define a time-dependent function called flow $\psi_{t}({\bm{x}}):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, which is diffeomorphism and satisfies $\psi_{t}(X_{1})=X_{t}$, to describe a continuous transform between $X_{1}$ and $X_{0}$. Specifically, for $0\leq s<t\leq 1$, we can sequentially transfer the sample $X_{t}$ to $X_{s}=\psi_{s}(X_{1})=\psi_{s}(\psi_{t}^{-1}(X_{t}))=\psi_{s|t}(X_{t})$. Here, the $\psi_{t}$ is uniquely characterized by a flow ODE
@@
Unfortunately, we cannot access ${\bm{v}}_{t}({\bm{x}}_{t})$ due to intractable integration over all ${\bm{x}}_{0}$. To address this, Lipman et al. (2023) proposes conditional flow matching:
@@
where the last equality is due to the definition of ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})$. One of the most important contribution of Lipman et al. (2023) is that the trained velocity field ${\bm{v}}_{t}^{\theta}({\bm{x}}_{t})$ from the conditional flow matching Eq. (3) can also approximate the unconditional velocity field ${\bm{v}}_{t}({\bm{x}}_{t})$ in Eq. (2). Accordingly, after training of the velocity field using Eq. (3), we can generate a sample by solving the following ODE:
@@
Without loss of generality, the problem could be extended to a flow defined in latent space. Accordingly, we will use ${\bm{x}}_{t}$ to denote both images and latent codes.
Among various flows, the affine conditional flow is defined as $\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})=a_{t}{\bm{x}}_{0}+b_{t}{\bm{x}}_{1}$ where $a_{0}=b_{1}=1$ and $a_{1}=b_{0}=0$, and the conditional velocity field derived from the rectified flow for Eq. (3) is
@@
In case of linear conditional flow (or rectified flow) Liu et al. (2023b), we set $a_{t}=t$ and $b_{t}=1-t$, and the conditional velocity field derived from the rectified flow for Eq. (3) is ${\bm{v}}_{t}({\bm{x}}_{t}|{\bm{x}}_{0})={\bm{x}}_{1}-{\bm{x}}_{0}$.
Figure 2: Overview. (a) Starting from the inverted latent, the ODE from source to target images can be obtained. (b) In contrast, existing inversion-free approaches suffer from nonsmooth trajectories, as ${\bm{q}}_{t}$ is sampled with random noise at each step, often resulting in editing artifacts. (c) FlowAlign uses the regularized velocity ${\bm{v}}_{t}^{{\bm{x}}}$ from the flow-matching loss as in Eq. (24), producing smoother and more consistent trajectories between the source and target images. Prompt: "…white and red sign that reads CAFE" $\rightarrow$ "…white and red sign that reads NeurIPS".
@@ 3.1 Text-based Image Editing using Pre-trained Flow Models
In text-based image editing using flow models, we aim to translate a source image ${\bm{x}}_{src}$ to a target image ${\bm{x}}_{tgt}$ based on text description of each image or editing instruction. In particular, such translation can be represented through a linear conditional flow between two image distributions,
$\psi_{t}({\bm{x}}_{src}|{\bm{x}}_{tgt}):={\bm{x}}_{t}=(1-t){\bm{x}}_{tgt}+t{\bm{x}}_{src},$ (7)
Note that the associated flow ODE described by Eq. (5) requires training between two image distributions using the conditional flow matching Eq. (3) with Eq. (7). To bypass the additional training, consider two generative flows that transfers the same noise ${\bm{\epsilon}}\sim\mathcal{N}(0,{\mathbf{I}}_{d})$ to each image,
@@
Importantly, the system of equations (Eq. (7), 8 and 9) lead to the following key equality:
@@ 3.2 Trajectory Errors from Inversion-Free Approaches
In Eq. (13), ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$ are first computed at each discrete time step $t$. If an inverted latent ${\bm{\epsilon}}_{\text{inv}}$ exists, obtained via ODE inversion from the source image ${\bm{x}}_{\text{src}}$ along the trajectory of ${\bm{q}}_{t}$ (as illustrated in Figure 2(a)), then ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$ can be further updated as:
@@
with the initialization condition ${\bm{p}}_{1}={\bm{q}}_{1}={\bm{\epsilon}}_{inv}$. However, avoiding the ODE inversion process is more advantageous, as it effectively reduces the overall computation by half. Toward this aim, FlowEdit Kulikov et al. (2024) proposed the following update for ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$
@@
However, one of the critical limitations of Eq. (15) is that ${\bm{q}}_{t}$ trajectory is not sufficiently smooth due to the random sampling of ${\bm{\epsilon}}$ (Figure 2(b)). This leads to the inaccuracy of using simulated ODE Eq. (13) for ${\bm{x}}_{t}$. To mitigate this, we propose an explicit trajectory regularization method in the following section.
3.3 Trajectory Regularization using Flow-Matching Loss
Given that the trajectory of ${\bm{x}}_{t}$ in Eq. (13) is not necessarily smooth due to the random sampling of ${\bm{\epsilon}}$, our objective is to simulate an ODE that explicitly penalizes deviations from a smooth trajectory as a form of regularization. Inspired by the recent advances of the optimal control approaches for flow models Rout et al. (2025a, b), we consider the following optimal control problem
@@ 3.3 Trajectory Regularization using Flow-Matching Loss
where
$\ell({\bm{x}}_{t},{\bm{u}}_{t},t)$ $:=\frac{1}{2}\|{\bm{u}}_{t}-\left(v_{t}^{\theta}({\bm{p}}_{t},c_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},c_{src})\right)\|^{2}+\frac{\lambda_{t}}{2}\|{\bm{u}}_{t}-({\bm{x}}_{tgt}-{\bm{x}}_{src})\|^{2}$ (18)
$m({\bm{x}}_{0})$ $:=\frac{\eta}{2}\|{\bm{x}}_{0}-{\bm{x}}_{tgt}\|^{2}$ (19)
with the regularization parameter $\lambda_{t},\eta>0$. Note that the first term in Eq. (18) enforces the similarity to the original velocity field in Eq. (13), whereas the second term is originated from the flow matching for the linear conditional flow:
$\mathcal{L}_{CFM}$ $=\mathbb{E}_{{\bm{x}}_{tgt}\sim q}\left[\mathcal{L}_{CFM,{\bm{x}}_{tgt}}\right]$ (20)
$\mathcal{L}_{CFM,{\bm{x}}_{tgt}}$ $=\mathbb{E}_{t\in[0,1]}\|{\bm{v}}_{t}({\bm{x}}_{t}|{\bm{x}}_{tgt})-{\bm{u}}_{t}\|^{2}=\int_{0}^{1}\|{\bm{x}}_{tgt}-{\bm{x}}_{src}-{\bm{u}}_{t}\|^{2}p(t)dt$ (21)
[…]
@@
$d{\bm{x}}_{t}$ $={\bm{v}}_{t}^{\bm{x}}({\bm{p}}_{t},{\bm{q}}_{t},{\bm{p}}_{0},{\bm{q}}_{0})dt,\quad{\bm{x}}_{1}={\bm{x}}_{src}$ (22)
${\bm{v}}_{t}^{\bm{x}}({\bm{p}}_{t},{\bm{q}}_{t},{\bm{p}}_{0},{\bm{q}}_{0})$ $:={\bm{v}}_{t}({\bm{p}}_{t},c_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},c_{src})+\frac{\lambda_{t}}{(1+\lambda_{t})t}\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (23)
as $\eta\rightarrow 0$, where $\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]={\bm{q}}_{t}-t{\bm{v}}_{t}({\bm{q}}_{t},c_{src}),\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]={\bm{p}}_{t}-t{\bm{v}}_{t}({\bm{p}}_{t},c_{tgt})$ are Tweedie’s denoising estimates.
@@
where we use a time-varying coefficient $\lambda_{t}=\frac{\gamma t}{1-\gamma t}$ to arrive at $\frac{\lambda_{t}}{(1+\lambda_{t})t}=\gamma$. This formulation ensures that $\lambda_{t}$ monotonically decreases toward zero as $t$ decreases from 1 to 0, implying that the flow-matching guidance is emphasized during the early stages of reverse sampling and gradually diminishes as the image becomes more refined.
The final velocity in Eq. (24) needs more discussion. Similar to the drift term in Eq. (13), the first term in Eq. (24) serves as the directional signal from the source to the target semantics. On the other hand, unlike the original flow matching penalty in Eq. (18), there is no dependency on ${\bm{x}}_{tgt}$ in the the second term in Eq. (24). Instead, it becomes a source consistent regularization gradient based on the distance between the clean estimates of ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$, computed using Tweedie formula. This regularization gradient keeps the trajectory ${\bm{p}}_{t}$ close to the source-consistency direction represented by $\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]$, thereby implicitly regulating ${\bm{x}}_{t}$. Finally, the assumption $\eta\rightarrow 0$ is to deal with the lack of paired ${\bm{x}}_{tgt}$ as the terminal condition.  [^2] One could also bypass this by assuming that ${\bm{x}}_{0}\rightarrow{\bm{x}}_{tgt}$ with a nonzero $\eta$. Alternatively, one could use the reward function of $R({\bm{x}})$ for the terminal condition, which is beyond the scope of this paper.
Another byproduct of trajectory regularization is the robustness to the classifier-free guidance (CFG). Unlike recent work Kulikov et al. (2024) that uses CFG to both estimated noises in with null-text embedding and different scales, we use the CFG as for the ODE trajectory ${\bm{p}}_{t}$ only,

### v3

@@ Abstract
Recent inversion-free, flow-based image editing methods such as FlowEdit leverages a pre-trained noise-to-image flow model such as Stable Diffusion 3, enabling text-driven manipulation by solving an ordinary differential equation (ODE). While the lack of exact latent inversion is a core advantage of these methods, it often results in unstable editing trajectories and poor source consistency. To address this limitation, we propose FlowAlign, a novel inversion-free flow-based framework for consistent image editing with optimal control-based trajectory control. Specifically, FlowAlign introduces source similarity at the terminal point as a regularization term to promote smoother and more consistent trajectories during the editing process. Notably, our terminal point regularization is shown to explicitly balance semantic alignment with the edit prompt and structural consistency with the source image along the trajectory. Furthermore, FlowAlign naturally supports reverse editing by simply reversing the ODE trajectory, highliting the reversible and consistent nature of the transformation. Extensive experiments demonstrate that FlowAlign outperforms existing methods in both source preservation and editing controllability.
@@ 1 Introduction
In text-based image editing meng2021sdedit ; mokady2023null ; tumanyan2023plug ; hertz2023delta ; kim2024dreamsampler ; kulikov2024flowedit ; brooks2023instructpix2pix , the goal is to transform a source image into a target based on either textual descriptions of the images or specific editing instructions. From a distributional perspective, the image editing task can be interpreted as a continuous normalizing flow (CNF) papamakarios2021normalizing that pushes forward a source distribution to a target distribution. Specifically, we parameterize a velocity field, that uniquely determines the flow, using a neural network. Then, the generative process corresponds to solving an ordinary differential equation (ODE) governed by the trained velocity field. To reduce the computational cost of simulating ODEs during training for likelihood evaluation, flow matching has been proposed lipman2023flow . Its conditional variant enables direct supervision by computing the target velocity, allowing efficient training of flow models between arbitrary distributions. These flow models includes score-based diffusion models song2020score ; ho2020denoising ; song2020denoising as well as rectified flow models liu2023flow ; esser2024scaling .
@@
Recently, within the framework of flow-based models, text-to-image generation has achieved significant advancements through improved time discretization, loss weighting, and model architecture, notably based on DiT peebles2023scalable . While these foundational models are trained to map samples from a normal distribution to a clean data distribution, additional training—either from scratch or via fine-tuning—is necessary to establish a flow between two arbitrary distributions.
@@
To mitigate this additional cost for constructing a new flow between image distributions, several approaches such as SDEdit meng2021sdedit and Dual Diffusion Implicit Bridge (DDIB) su2022dual have leveraged pre-trained noise-to-image diffusion models. However, SDEdit requires careful selection of an appropriate initial noise level for editing, while DDIB relies on an inversion process that is prone to errors arising from discretization and an approximated velocity field for subsequent timesteps.
Recently, RF-inversion rout2025semantic proposes an optimal-control based inversion for flow models, in which a guiding vector field steers the dynamics toward high-likelihood samples via velocity interpolation. Unfortunately, it still involves computational overhead due to the ODE inversion. To address this, FlowEdit kulikov2024flowedit proposed simulating an ODE between two image samples without inversion. However, empirical results showed that the method is quite sensitive to the hyperparameters and often fails to retain the source consistency. Moreover, the method heuristically applies classifier-free-guidance (CFG) to the both source and target velocity fields with different scaling factor and relies on skipping early timesteps of the ODE to enhance editing quality. These design choices introduce multiple hyperparameters to be searched and compromise the deterministic nature of ODE (see Section 1).
@@
Motivated by the observation that these limitations of FlowEdit arise from nonsmooth and unstable editing trajectories—partly due to the lack of explicit latent inversion—we introduce FlowAlign, an optimal control-based inversion-free approach for consistent and controllable text-driven image editing through trajectory regularization. In contrast to RF-inversion rout2025semantic that requires an ODE inversion, we introduces a structural similarity at the terminal point as our trajectory regularization term to overcome the instability caused by the absence of the inverted latent. Although the regularization is primarily enforced at the terminal point, we further observe that it also naturally enforces source consistency along the trajectory by penalizing unnecessary deviations from the original image.
@@
Our method is also computationally efficient, requiring only one additional function evaluation (NFE) per ODE step for the regularization term. In contrast, methods like FlowEdit kulikov2024flowedit incur twice the NFE due to the reliance on classifier-free guidance. Despite this efficiency, FlowAlign achieves superior source preservation and competitive or improved editing quality compared to existing approaches in image, video, and 3D editing. Finally, FlowAlign supports editing through backward ODE with the learned flow field, enabling accurate reconstruction of the original image from the edited output. This highlights the reversible and deterministic nature of the learned transformation, made possible by our explicit trajectory regularization.
@@ 2 Backgrounds
Suppose we have access to samples from two distributions $X_{1}\sim p$ and $X_{0}\sim q$ that forms independent coupling, $\pi_{0,1}(X_{0},X_{1})=p(X_{1})q(X_{0})$. We can define a time-dependent function called flow $\psi_{t}({\bm{x}}):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, which is diffeomorphism and satisfies $\psi_{t}(X_{1})=X_{t}$, to describe a continuous transform between $X_{1}$ and $X_{0}$. Specifically, for $0\leq s<t\leq 1$, we can sequentially transfer the sample $X_{t}$ to $X_{s}=\psi_{s}(X_{1})=\psi_{s}(\psi_{t}^{-1}(X_{t}))=\psi_{s|t}(X_{t})$. Here, the $\psi_{t}$ is uniquely characterized by a flow ODE
@@
Unfortunately, we cannot access ${\bm{v}}_{t}({\bm{x}}_{t})$ due to intractable integration over all ${\bm{x}}_{0}$. To address this, lipman2023flow proposes conditional flow matching:
@@
where the last equality is due to the definition of ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})$. One of the most important contribution of lipman2023flow is that the trained velocity field ${\bm{v}}_{t}^{\theta}({\bm{x}}_{t})$ from the conditional flow matching Eq. (3) can also approximate the unconditional velocity field ${\bm{v}}_{t}({\bm{x}}_{t})$ in Eq. (2). Accordingly, after training of the velocity field using Eq. (3), we can generate a sample by solving the following ODE:
@@
Among various flows, the affine conditional flow defined as $\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})=a_{t}{\bm{x}}_{0}+b_{t}{\bm{x}}_{1}$ is widely used, where $a_{0}=b_{1}=1$ and $a_{1}=b_{0}=0$. The resulting conditional velocity field derived from the rectified flow for Eq. (3) is then given by
@@
In case of linear conditional flow (or rectified flow) liu2023flow , $a_{t}=t$ and $b_{t}=1-t$, and the conditional velocity field derived from the rectified flow for Eq. (3) is ${\bm{v}}_{t}({\bm{x}}_{t}|{\bm{x}}_{0})={\bm{x}}_{1}-{\bm{x}}_{0}$.
Without loss of generality, the problem formulation could be extended to a flow defined in latent space. Accordingly, we will use ${\bm{x}}_{t}$ to denote both images and latent codes.
Figure 2: Overview. (a) Starting from the inverted latent, the ODE from source to target images can be obtained. (b) In contrast, existing inversion-free approaches suffer from nonsmooth trajectories, as ${\bm{q}}_{t}$ is sampled with random noise at each step, often resulting in editing artifacts. (c) FlowAlign uses the regularized velocity ${\bm{v}}_{t}^{{\bm{x}}}$ from the similarity regularization at the terminal point, producing smoother and more consistent trajectories between the source and target images. Prompt: "…white and red sign that reads CAFE" $\rightarrow$ "…white and red sign that reads NeurIPS".
@@ 3.1 Text-based Image Editing using Pre-trained Flow Models
In text-based image editing using flow models, we aim to translate a source image ${\bm{x}}_{src}$ at $t=1$ to a target image ${\bm{x}}_{tgt}$ at $t=0$ based on text description of each image or editing instruction. In particular, such translation can be represented through a linear conditional flow between two image distributions,
$\psi_{t}({\bm{x}}_{src}|{\bm{x}}_{tgt}):={\bm{x}}_{t}=(1-t){\bm{x}}_{tgt}+t{\bm{x}}_{src}\quad.$ (7)
Note that the associated flow ODE described by Eq. (5) requires training between two image distributions using the conditional flow matching of Eq. (3) with Eq. (7). To bypass the additional training, consider two generative flows that transfers the same noise ${\bm{\epsilon}}\sim\mathcal{N}(0,{\mathbf{I}}_{d})$ to each image,
@@
Importantly, the system of equations (Eq. (7), Eq. (8) and Eq. (9)) lead to the following key equality:
@@ 3.2 Trajectory Errors from Inversion-Free Approaches
In Eq. (13), ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$ are first computed at each discrete time step $t$. If an inverted latent ${\bm{\epsilon}}_{\text{inv}}$ is obtained via ODE inversion from the source image ${\bm{x}}_{\text{src}}$ along the trajectory of ${\bm{q}}_{t}$ (as illustrated in Figure 2(a)), then ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$ can be further updated as:
@@
with the initialization condition ${\bm{p}}_{1}={\bm{q}}_{1}={\bm{\epsilon}}_{inv}$. However, ODE inversion increase the overall computational cost. Thus, we are interested in avoiding the ODE inversion process. Toward this aim, FlowEdit kulikov2024flowedit proposed the following update for ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$
@@
Unfortunately, one of the critical limitations of FlowEdit using Eq. (15) is that ${\bm{q}}_{t}$ trajectory is not sufficiently smooth due to the random sampling of ${\bm{\epsilon}}$ (Figure 2(b)). This leads to the inaccuracy of using simulated ODE in Eq. (13) for ${\bm{x}}_{t}$. To mitigate this, we propose an explicit trajectory regularization method using the structural similarity at the terminal point.
3.3 Trajectory Regularization using Similarity at Terminal Point
Given that the trajectory of ${\bm{x}}_{t}$ in Eq. (13) is not necessarily smooth due to the random sampling of ${\bm{\epsilon}}$, our objective is to simulate an ODE that explicitly penalizes deviations from a smooth trajectory as a form of regularization. Inspired by the recent advances of the optimal control approaches for flow models rout2025semantic ; rout2025rbmodulation , we consider the following time-reversal optimal control problem
@@ 3.3 Trajectory Regularization using Similarity at Terminal Point
In this paper, we use the following loss:
$\ell({\bm{x}}_{t},{\bm{u}}_{t},t)$ $:=\frac{1}{2}\|{\bm{u}}_{t}-\left(v_{t}^{\theta}({\bm{p}}_{t},c_{tgt})-{\bm{v}}_{t}^{\theta}({\bm{q}}_{t},c_{src})\right)\|^{2},$ (18)
thereby enforcing the similarity of the optimal control ${\bm{u}}_{t}$ to the original velocity field in Eq. (13). Unfortunately, this is shown not sufficient for regularizing the trajectory deviation owing to the lack of inverted latent. Therefore, our goal is to utilize the terminal loss $m({\bm{x}}_{0})$ to enforce the trajectory smoothness.
One of the most important contributions of this work is showing that a similarity regularization at the terminal point serves the goal. Specifically, we introduce a $l_{2}$-based terminal point regularization:
${\bm{m}}({\bm{x}}_{0})=\frac{\eta}{2}\|{\bm{x}}_{0}-{\bm{x}}_{{src}}\|^{2}\quad$ (19)
This implies that the ODE evolution from from $t=1$ to $t=0$ ultimately converges to a terminal solution that closely resembles the starting point, i.e. ${\bm{x}}_{tgt}\simeq{\bm{x}}_{src}$. While this might raise concerns about convergence to a trivial solution where ${\bm{x}}_{t}={\bm{x}}_{src}$ for all $t$, the crucial distinction lies in our use of a finite $\eta$, as opposed to taking the limit $\eta\rightarrow\infty$ as done in RF-inversion rout2025semantic . In what follows, we show that this terminal point regularization lead to the balance between the semantic guidance and structural preservation.
@@
$d{\bm{x}}_{t}$ $={\bm{v}}_{t}^{\bm{x}}({\bm{p}}_{t},{\bm{q}}_{t},{\bm{p}}_{0},{\bm{q}}_{0})dt$ (20)
with the initial condition ${\bm{x}}_{1}={\bm{x}}_{src}$, where
${\bm{v}}_{t}^{\bm{x}}({\bm{p}}_{t},{\bm{q}}_{t},{\bm{p}}_{0},{\bm{q}}_{0})$ $:\simeq{\bm{v}}_{t}({\bm{p}}_{t},c_{tgt})-{\bm{v}}_{t}({\bm{q}}_{t},c_{src})+\gamma\left(\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]-\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]\right)$ (21)
where $\gamma=\frac{-\eta}{1-\eta t}$ is positive for sufficiently large $\eta$ and $\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]={\bm{q}}_{t}-t{\bm{v}}_{t}({\bm{q}}_{t},c_{src}),\mathbb{E}[{\bm{p}}_{0}|{\bm{p}}_{t}]={\bm{p}}_{t}-t{\bm{v}}_{t}({\bm{p}}_{t},c_{tgt})$ are Tweedie’s denoising estimates.
@@
The final velocity in Eq. (22) needs more discussion. Similar to the drift term in Eq. (13), the first term in Eq. (22) serves as the directional signal from the source to the target semantics. On the other hand, the second term becomes a source consistent regularization gradient based on the distance between the clean estimates of ${\bm{p}}_{t}$ and ${\bm{q}}_{t}$, computed using Tweedie formula. This regularization gradient keeps the trajectory ${\bm{p}}_{t}$ close to the source-consistency direction represented by $\mathbb{E}[{\bm{q}}_{0}|{\bm{q}}_{t}]$, thereby implicitly regulating ${\bm{x}}_{t}$.
Another byproduct of trajectory regularization is the robustness to the classifier-free guidance (CFG). Unlike recent work kulikov2024flowedit that uses CFG to both estimated noises in with null-text embedding and different scales, we use the CFG as for the ODE trajectory ${\bm{p}}_{t}$ only,